MICROWAVE SPECTRA AND STRUCTURES OF H_{4}C_{2}⋯AgCl AND H_{4}C_{2}⋯CuCl

\underline{N. R. WALKER}, S. L. STEPHENS, V. A. MIKHAILOV AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

A Balle-Flygare FT-MW spectrometer coupled to a laser ablation source has been used to measure the pure rotational spectra of H_{4}C_{2}⋯AgCl and H_{4}C_{2}⋯CuCl. Both molecules are generated via laser ablation (532 nm) of a metal rod in the presence of CCl_{4}, C_{2}H_{4} and argon and are stabilized by supersonic expansion. Rotational constants (A_{0}, B_{0}, C_{0}) and the centrifugal distortion constant, D_{J} , have been measured for six isotopologues of H_{4}C_{2}⋯AgCl and five isotopologues of H_{4}C_{2}⋯CuCl with substitutions at the metal, chlorine and carbon atoms in each case. The spectrum of each molecule is consistent with a C_{2v} structure in which the metal atom interacts with the π-orbital on ethene. The measured rotational constants allow determination of the length of the bond between the metal and chlorine atoms, r_{MCl}, and the distance between the metal atom and the centre of the ethene double bond, r_{MEt}. Nuclear quadrupole coupling constants have been determined for the chlorine atom in each molecule and also for copper in H_{4}C_{2}⋯CuCl.